OpenStudy (anonymous):

Solve for x. Show your work √x-7 +5=11

OpenStudy (abb0t):

is it \(\sqrt{x-7}\) or \(\sqrt{x}-7\) ??

OpenStudy (anonymous):

\[\sqrt{x-7}\]

OpenStudy (abb0t):

1. Subtract 5 to both sides. 2. square both sides 3. add 7 to both sides.

OpenStudy (anonymous):

x^2=22^2?